# On an Optimal Stopping Problem with a Discontinuous Reward

### v3

@@ Abstract.
Keywords: Variable annuities, Optimal stopping, Surrender option, Discontinuous reward, Free-boundary problems
@@ 1. Introduction
Variable annuities (VA) are structured products sold by insurance companies that are mainly used for retirement planning. An initial premium is deposited in an investment account (or fund) whose return is linked to that of one or more risky assets. They are similar to mutual funds, but they also offer financial guarantees at the end of a pre-determined accumulation period. The embedded guarantees are akin to long-dated options, but they are funded by a periodic fee, generally set as a percentage of the investment account value, rather than being paid upfront. In this paper, we focus on a guaranteed minimum maturity benefit (GMMB), which provides the policyholder a minimum guaranteed amount at maturity of the contract. Another common feature of variable annuities contracts is that policyholders generally have the right to surrender, or lapse, their contract prior to maturity. When policyholders choose to do so, they receive the value accumulated in the investment account, reduced by a penalty for early surrender (or surrender charge). The uncertainty faced by VA providers with respect to early termination is known as surrender risk. Early surrenders can entail negative consequences such as liquidity issues. Thus, incorporating adequate surrender assumptions in the pricing of VAs is essential for risk management purposes; see Niittuinperä (2022).
Different early surrender modelling approaches have been proposed in the literature (see Bauer et al. (2017) and Feng et al. (2022) for a review), from the use of utility functions, Gao and Ulm (2012), to modern statistical techniques, Zhu and Welsch (2015). Early surrenders are also often expressed as a decision taken by policyholders on a strictly rational basis, meaning that the contract is terminated as soon as it is optimal to do so from a financial perspective (see Bernard et al. (2014a), Jeon and Kwak (2018), Kang and Ziveyi (2018), MacKay et al. (2023), and Milevsky and Salisbury (2001), among others). Bauer et al. (2017) discusses the impact of other factors on policyholder behavior, such as taxes and expenses. This article gave rise to another stream of literature in which market frictions such as taxation rules are considered (Alonso-García et al. (2022), Bauer and Moenig (2023), and Moenig and Bauer (2016)), which helps to explain the discrepancies between market and model fee rates. Other recent studies include lapse and reentry strategies in their analysis (Bernard and Moenig (2019) and Moenig and Zhu (2018)), whereas Moenig and Zhu (2021) consider a third-party investor to whom the policyholder can sell her contract.
@@
The time discontinuity and the unboundedness of the reward function involved in the pricing of variable annuities with GMMB prevent the simple application of results from the American option pricing literature to our problem. For example, to express the value function as the solution to a free-boundary value problem, one needs to establish its continuity. However, continuity of the value function often follows from the continuity of the reward function, which, in our setting, is unbounded and only upper semi-continuous, making the results usually cited in the context of American options inapplicable (see, for instance, Bassan and Ceci (2002), Bensoussan and Lions (1982), De Angelis and Stabile (2019), Jaillet et al. (1990), Krylov (1980), and Lamberton (2009)). In this paper, following the work of Van Moerbeke (1974) and Palczewski and Stettner (2010) (in the context of impulse control problems), we establish continuity of the value function by showing that it admits an alternative representation in terms of a continuous reward function. We can then confirm that it is a solution to a free-boundary value problem and prove further properties, which provides enough regularity to apply a generalized version of Itô’s formula to the value process. This allows us to derive various integral representations for the value function, thus extending the results of Bernard et al. (2014b) to general time-dependent fee and surrender charge functions and introducing a new decomposition, the continuation premium representation, to the literature on VA pricing.
The shape and the (non-)emptiness of the exercise region of an American option with a time-homogeneous reward function has first been studied rigorously by Villeneuve (1999). Other authors, such as Jönsson et al. (2006), Kotlow (1973), and Villeneuve (2007), also studied the conditions under which the exercise region has a particular shape. These results are however not directly applicable to the problem we study in this paper, because of the time-dependence and the discontinuity of our reward function. Under the assumption of a general fee and surrender charge structure, we are nonetheless able to identify a condition, expressed as a partial differential inequality, under which the optimal stopping time is always at maturity of the contract, thus mitigating surrender risk. This condition provides a powerful tool to understand the interplay between the fees and surrender charges, and its effect on the optimal stopping strategy. It can also be used to further study and characterize the shape of the optimal exercise, or surrender, region. In particular, we show that the surrender region can be a disconnected set and that the optimal surrender boundary can be discontinuous, thus illustrating a major difference between the pricing of VA contracts and standard American options. Our results also extend the work of Milevsky and Salisbury (2001), who consider an infinite horizon, to a more general finite horizon problem.
@@
  • We present an alternative representation for the value function of the VA contract in terms of a continuous reward function. As stated above, various results in the literature, in particular those regarding continuity of the value function, rely on the continuity of the reward function. In our paper, this alternative representation is used in Section 4 to perform a rigorous theoretical study of the value function of the VA contract and to show that it solves a free boundary value problem. These results justify the use of different numerical methods already applied to VA pricing in the actuarial literature.
  • We provide two further representations for the value function. For the first one, we decompose the value of the contract into the value of the maturity benefit and an integral term akin to the early exercise premium from the American option literature. The second representation is new to the actuarial and American option pricing literature and expresses the contract value as the sum of the surrender benefit and an integral term coined the continuation premium. These two representations are used in Section 4.3 to better understand the value function of a VA contract and in the proofs of Section 4.4, which presents characterizations of the (non-) emptiness of the surrender region. The two representations can also be used to develop numerical pricing algorithms.
@@ 2.1. Market Model
On a probability space $(\Omega,\mathcal{F},\mathbb{Q})$, let $W=\{W_{t}\}_{t\geq 0}$ be a standard Brownian motion whose augmented filtration is denoted by $\mathbb{F}=\{\mathcal{F}_{t}\}_{t\geq 0}$. $\mathbb{Q}$ is the probability measure used to price the assets presented below.
@@
where $r$, $\sigma$ are deterministic constants with $\sigma>0$. It is easy to verify that (1) has a unique strong solution given by $B_{t}=B_{0}e^{rt}$ and
$S_{t}=S_{0}e^{(r-\frac{\sigma^{2}}{2})t+\sigma W_{t}}.$
@@ 2.2. Variable Annuity Contract
We consider a simplified variable annuity contract offering a guaranteed minimum accumulation benefit at maturity $T\in\mathbb{R}_{+}$, with $\mathbb{R}_{+}$ denoting the strictly positive real numbers, that is +=$\mathbb{R}_{+}=\{x\in\mathbb{R}|x>0\}$. At inception of the contract, the policyholder deposits an initial premium in an investment account (or fund) tracking the financial market. Here, we assume that this account tracks the risky asset with price process $S$. We denote by $F=\{F_{t}\}_{0\leq t\leq T}$ the value process of the account.
We remark that variable annuity policyholders usually benefit from a return-of-premium guarantee if they die before maturity. Under the assumption that mortality risk is completely diversifiable, it is straight-forward to add the death benefit in the analysis of the contract, see for example MacKay et al. (2017). See also De Angelis et al. (2024) for a more in-depth analysis. In this paper, we ignore mortality risk.
@@
Should she decide to surrender her contract prior to maturity, the policyholder receives the amount accumulated in the account subject to a penalty, expressed as a percentage of the account, which can depend on time and the value of the account. If no surrenders occur, the maturity benefit is paid at $T$.
@@
where $g:[0,\,T]\times\mathbb{R}_{+}\rightarrow(0,1]$ is in $C^{1,2}$ with bounded derivatives on $[0,T]\times\mathbb{R}_{+}$, is non-decreasing in $t$ and satisfies $g(T,x)=1$ for all $x\in\mathbb{R}_{+}$. This function represents the amount received by the policyholder upon surrender at $t<T$ or at maturity $T$, given that the account has value $x$.
In the literature, $1-g$ is often called the surrender charge. It is non-increasing in time; a later surrender will yield a higher proportion of the account value. In this paper, we use the term surrender charge function to refer to either $g$ or $1-g$.
@@
for $(t,x)\in[0,T]\times\mathbb{R}_{+}$, where $\mathcal{T}_{t,\,T}$ is the set of all stopping times taking value in the interval $[t,\,T]$. Moreover, throughout the paper, it must be understood that for $s>t$,
$v(s,F^{t,x}_{s})=\operatorname*{ess\,sup}_{\tau\in\mathcal{T}_{s,\,T}}\mathbb{E}[e^{-r(\tau-s)}\varphi(\tau,F_{\tau}^{t,\,x})],$
@@
The difference between the value of the full variable annuity contract and the present value of the maturity benefit is the value of the surrender right, which we denote by $e:[0,T]\times\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}\cup\{0\}$ and define as
$e(t,x):=v(t,x)-h(t,x).$
@@ 3. Optimal Stopping Time
For a general reward function $\varphi$, such an optimal stopping time does not necessarily exist. However, under some regularity conditions on the discounted reward process ${Z=\{Z_{t}\}_{0\leq t\leq T}}$, with $Z_{t}:=e^{-rt}\varphi(t,F_{t})$, the existence and the form of an optimal stopping time are well known from the theory of optimal stopping for random processes in continuous time, see, for instance, Karatzas and Shreve (1998) (Appendix D, Theorem D.12), Peskir and Shiryaev (2006) (Theorem 2.2), Lamberton (1998) (Section 10.2.1), Lamberton (2009) (Theorem 2.3.5 and Section 2.3.4), El Karoui (1981), Chapter 2. However, with the exception of El Karoui (1981), these results rely on the (almost sure) continuity of the reward process, which does not hold for the one involved in variable annuity pricing since its trajectories are discontinuous at time $T$ with positive probability. In fact the trajectories of $Z$ are upper semi-continuous. Thus, additional work is required to show the existence of an optimal stopping time in our setting.
3.1. Surrender Right - Trivial Case
@@
We first observe that for $0\leq t<T$, $Y_{t}=Z_{t}$ so that if $Y$ is a submartinale, then for any $0\leq s\leq t<T$, $\mathbb{E}_{s}[Z_{t}]\geq Z_{s}$.
For $0\leq s<t=T$, since $g(T,x)=1$ for all $x\in\mathbb{R}_{+}$, we have
@@
An application of Itô’s lemma to the discounted surrender value process $Y=\{Y_{t}\}_{0\leq t\leq T}$, with $Y_{t}:=e^{-rt}F_{t}g(t,\,F_{t})$, yields
@@
for all $(t,x)\in[0,T)\times\mathbb{R}_{+}$, $Y$ is a submartingale and so is the discounted reward process by Lemma 3.1. The optimal stopping time follows from Lemma 3.2. ∎
@@
Next, we present some applications of Proposition 3.1.
@@
Let the fee rate be constant, that is $C(t,x)=c>0$, and let ${g(t,x)=e^{-k(T-t)}}$ for some $k>0$. To eliminate the incentive to surrender, it suffices to choose $k$ and $c$ such that (9) holds; a quick calculation yields $k\geq c$. In this setting, it suffices to set the surrender rate $k$ at or above the fee rate $c$ to eliminate optimal early surrenders. This result is well-known and discussed in Bernard and MacKay (2015), Proposition 3.1 and MacKay (2014), Proposition 4.4.2. In particular, MacKay (2014) shows that in the Black-Scholes setting, $g(t,x)=e^{-c(T-t)}$ is the maximal surrender charge function that can be used to remove the surrender incentive when the fee rate is constant.
@@
Let $g(t,x)=1-k\left(1-\frac{t}{T}\right)^{3}$ for some constant $k>0$, as in Bacinello and Zoccolan (2019), MacKay (2014) Section 4.3.2., and MacKay et al. (2017). As explained in MacKay (2014), this form of surrender charge function mimics the surrender charges in the market, which are usually high in the first years of the contract and drop drastically thereafter. Assume also that the fee function depends only on time such that $C(t,\cdot)=c(t)$. In this example, we do not specify a particular form for the fee function; (9) is used to obtain the form $c(t)$ eliminating the surrender incentive. Simple calculations show that (9) is satisfied if
@@
Thus, any function $c(t)$ that satisfies this inequality above will eliminate the surrender incentive (from a risk-neutral perspective). We may simply set $c(t)=\frac{3k/T\left(1-t/T\right)^{2}}{1-k(1-t/T)^{3}}$. This example highlights the interplay between the fee and the surrender charge structures. In this setting, it would have been impossible to satisfy (9) using a constant fee rate, except with $c=0$, as the function on the right-hand side of (10) vanishes at maturity. Having a zero fee rate throughout the life of the VA contract is not a feasible solution from the insurer’s perspective since the maturity guarantee must be financed. On the other hand, isolating $k$ in (10) gives
@@
The examples presented above illustrate the interplay between fees, surrender charges, and surrender incentives. In order to eliminate early surrender incentives, surrender charges and fees must be structured conjointly; that is, a time-dependent fee must be paired with a time-dependent surrender charge to satisfy (9).
@@
3.2. Surrender Right - Non-Trivial Case
In this section, we study the existence of an optimal stopping time for the variable annuity contract in the general case, including when condition (9) is not satisfied. We first recall two definitions from probability and optimal stopping theory.

### v4

@@ Abstract.
Keywords: Variable annuities, Optimal stopping, Surrender option, Discontinuous reward, Free boundary problems
@@ 1. Introduction
Variable annuities (VA) are structured products sold by insurance companies that are mainly used for retirement planning. An initial premium is deposited in an investment account (or fund) whose return is linked to that of one or more risky assets. They are similar to mutual funds, but they also offer financial guarantees at the end of a pre-determined accumulation period. The embedded guarantees are akin to long-dated options, but they are funded by a periodic fee, generally set as a percentage of the investment account value, rather than being paid upfront. In this paper, we focus on a guaranteed minimum maturity benefit (GMMB), which provides the policyholder a minimum guaranteed amount at maturity of the contract. Another common feature of variable annuity contracts is that policyholders generally have the right to surrender, or lapse, their contract prior to maturity. When policyholders choose to do so, they receive the value accumulated in the investment account, reduced by a penalty for early surrender (or surrender charge). The uncertainty faced by VA providers with respect to early termination is known as surrender risk. Early surrenders can entail negative consequences such as liquidity issues. Thus, incorporating adequate surrender assumptions in the pricing of VAs is essential for risk management purposes; see Niittuinperä (2022).
Different early surrender modelling approaches have been proposed in the literature (see Bauer et al. (2017) and Feng et al. (2022) for a review), from the use of utility functions, Gao and Ulm (2012), to modern statistical techniques, Zhu and Welsch (2015). Early surrenders are also often expressed as a decision taken by policyholders on a strictly rational basis, meaning that the contract is terminated as soon as it is optimal to do so from a risk-neutral perspective (see Bernard et al. (2014a), Jeon and Kwak (2018), Kang and Ziveyi (2018), MacKay et al. (2023), and Milevsky and Salisbury (2001), among others). Bauer et al. (2017) discusses the impact of other factors on policyholder behavior, such as taxes and expenses. This article gave rise to another stream of literature in which market frictions such as taxation rules are considered (Alonso-García et al. (2022), Bauer and Moenig (2023), and Moenig and Bauer (2016)), which helps to explain the discrepancies between market and model fee rates. Other recent studies include lapse and reentry strategies in their analysis (Bernard and Moenig (2019) and Moenig and Zhu (2018)), whereas Moenig and Zhu (2021) consider a third-party investor to whom the policyholder can sell her contract.
@@
The time discontinuity and the unboundedness of the reward function involved in the pricing of variable annuities with GMMB prevent the simple application of results from the American option pricing literature to our problem. For example, to express the value function as the solution to a free boundary value problem, one needs to establish its continuity. However, continuity of the value function often follows from the continuity of the reward function, which, in our setting, is unbounded and only upper semi-continuous, making the results usually cited in the context of American options inapplicable (see, for instance, Bassan and Ceci (2002), Bensoussan and Lions (1982), De Angelis and Stabile (2019), Jaillet et al. (1990), Krylov (1980), and Lamberton (2009)). In this paper, following the work of Van Moerbeke (1974) and Palczewski and Stettner (2010) (in the context of impulse control problems), we establish continuity of the value function by showing that it admits an alternative representation in terms of a continuous reward function. We can then confirm that it is a solution to a free boundary value problem and prove further properties, which provides enough regularity to apply a generalized version of Itô’s formula to the value process. This allows us to derive various integral representations for the value function, thus extending the results of Bernard et al. (2014b) to general time-dependent fee and surrender charge functions and introducing a new decomposition, the continuation premium representation, to the literature on VA pricing.
The shape and the (non-)emptiness of the exercise region of an American option with a time-homogeneous reward function has first been studied rigorously by Villeneuve (1999). Other authors, such as Jönsson et al. (2006), Kotlow (1973), and Villeneuve (2007), also studied the conditions under which the exercise region has a specific shape. These results are however not directly applicable to the problem we study in this paper, because of the time-dependence and the discontinuity of our reward function. Under the assumption of a general fee and surrender charge structure, we are nonetheless able to identify a condition, expressed as a partial differential inequality, under which the optimal stopping strategy is always to hold on to the contract until maturity, thus mitigating surrender risk. This condition provides a powerful tool to understand the interplay between the fees and surrender charges, and its effect on the optimal stopping strategy. It can also be used to further study and characterize the shape of the optimal exercise, or surrender, region. In particular, we show that the surrender region can be a disconnected set and that the optimal surrender boundary can be discontinuous, thus further highlighting the complexity of this optimal stopping problem. We also extend the work of Milevsky and Salisbury (2001), who consider an infinite horizon, to a more general finite horizon problem.
@@
  • We present an alternative representation for the value function of the VA contract in terms of a continuous reward function. As stated above, this representation makes it easy to show continuity of the value function. In our paper, this alternative representation is used in Section 4 to show that the value function solves a free boundary value problem. These results justify the use of different numerical methods already applied to VA pricing in the actuarial literature.
  • We provide two further representations for the value function. For the first one, we decompose the value of the contract into the value of the maturity benefit and an integral term akin to the early exercise premium from the American option literature. The second representation is new to the actuarial and American option pricing literature and expresses the contract value as the sum of the surrender benefit and an integral term coined the continuation premium. These two representations are used in Section 4.4 to characterize the (non-) emptiness of the surrender region. The two representations can also be used to develop numerical pricing algorithms.
@@ 2.1. Market Model
On a probability space $(\Omega,\mathcal{F},\mathbb{Q})$, let $W=\{W_{t}\}_{t\geq 0}$ be a standard Brownian motion whose augmented filtration is denoted by $\mathbb{F}=\{\mathcal{F}_{t}\}_{t\geq 0}$. $\mathbb{Q}$ is the pricing (probability) measure in the market presented below.
@@
where $r$, $\sigma$ are deterministic constants with $r\geq 0$ and $\sigma>0$. It is easy to verify that (1) has a unique strong solution given by $B_{t}=B_{0}e^{rt}$ and
$S_{t}=S_{0}e^{\left(r-\frac{\sigma^{2}}{2}\right)t+\sigma W_{t}}.$
@@ 2.2. Variable Annuity Contract
We consider a simplified variable annuity contract offering a guaranteed minimum accumulation benefit at maturity $T\in\mathbb{R}_{+}$, with $\mathbb{R}_{+}=\{x\in\mathbb{R}|x>0\}$. At inception of the contract, the policyholder deposits an initial premium in an investment account (or fund) tracking the risky asset. We denote by $F=\{F_{t}\}_{0\leq t\leq T}$ the value process of the account.
We remark that variable annuity policyholders usually benefit from a return-of-premium guarantee if they die before maturity. Under the assumption that mortality risk is completely diversifiable, it is straight-forward to add the death benefit in the analysis of the contract, see for example MacKay et al. (2017) and De Angelis et al. (2024). In this paper, we ignore mortality risk.
@@
Should she decide to surrender her contract prior to maturity, the policyholder receives the amount accumulated in the account subject to a penalty, expressed as a percentage that can depend on time and the value of the account. If no surrenders occur, the maturity benefit is paid at $T$.
@@
where $g:[0,\,T]\times\mathbb{R}_{+}\rightarrow(0,1]$ is in $C^{1,2}$ with bounded derivatives on $0,T]\times\mathbb{R}_{+}$, is non-decreasing in $t$ and satisfies $g(T,x)=1$ for all $x\in\mathbb{R}_{+}$. Thus, $\varphi(t,x)$ represents the amount received by the policyholder upon surrender at $t<T$ or at maturity $T$, given that the account has value $x$.
In the literature, 1g1-g is often called the surrender charge. It is non-increasing in time; a later surrender will yield a higher proportion of the account value. In this paper, we use the term surrender charge function to refer to either $g$ or 1g1-g.
@@
for $(t,x)\in[0,T]\times\mathbb{R}_{+}$, where $\mathcal{T}_{t,\,T}$ is the set of all stopping times taking value in the interval [t,T][t,\,T]. Moreover, throughout the paper, it must be understood that for $s>t$,
$v(s,F^{t,x}_{s})=\operatorname*{ess\,sup}_{\tau\in\mathcal{T}_{s,\,T}}\mathbb{E}[e^{-r(\tau-s)}\varphi(\tau,F_{\tau}^{t,\,x})|\mathcal{F}_{s}],$
@@ 3. Optimal Stopping Time
For a general reward function $\varphi$, such an optimal stopping time does not necessarily exist. However, under some regularity conditions on the discounted reward process $Z=\{Z_{t}\}_{0\leq t\leq T}$}, with $Z_{t}:=e^{-rt}\varphi(t,F_{t})$, the existence and the form of an optimal stopping time are well known from the theory of optimal stopping for random processes in continuous time, see, for instance, Karatzas and Shreve (1998) (Appendix D, Theorem D.12), Peskir and Shiryaev (2006) (Theorem 2.2), Lamberton (1998) (Section 10.2.1), Lamberton (2009) (Theorem 2.3.5 and Section 2.3.4), El Karoui (1981), Chapter 2. With the exception of El Karoui (1981), these results rely on the (almost sure) continuity of the reward process, which does not hold for the one involved in variable annuity pricing since its trajectories are discontinuous at time $T$ with positive probability. In fact the trajectories of $Z$ are upper semi-continuous. Thus, additional work is required to show the existence of an optimal stopping time in our setting.
3.1. Trivial Case
@@
We first observe that for $0\leq t<T$, $Y_{t}=Z_{t}$ so that if $Y$ is a submartingale, then for any $0\leq s\leq t<T$, $\mathbb{E}_{s}[Z_{t}]\geq Z_{s}$. For $0\leq s<t=T$, since $g(T,x)=1$ for all $x\in\mathbb{R}_{+}$, we have
@@
An application of Itô’s lemma to the discounted surrender value process $Y$, defined in Lemma 3.1, yields
@@
for all $(t,x)\in[0,T)\times\mathbb{R}_{+}$, $Y$ is a local submartingale. Furthermore, since $Y_{t}\leq e^{-rt}F_{t}\leq e^{-rt}S_{t}$, it follows from the properties of the geometric Brownian motion that for any $t<\infty$,
$\sup_{\tau\leq t}\mathbb{E}[Y_{\tau}^{2}]\leq\sup_{\tau\leq t}\mathbb{E}[(e^{-r\tau}S_{\tau})^{2}]<\infty,$
showing that $\{Y_{\tau}\}_{\tau\leq t}$ is uniformly integrable for any fixed $t<\infty$, and thus of class (DL) (see Definition 3.2). This is sufficient to show that $Y$ is a true submartingale (see Proposition 1.7 of Chapter IV of Revuz and Yor (1999)). Then, by Lemma 3.1, the discounted reward process $\{Z_{t}\}_{0\leq t\leq T}$} is also a submartingale, and the optimal stopping time follows from Lemma 3.2. ∎
@@
To give a better idea of the impact of the fee and the surrender charge functions on the surrender incentives, we present some applications of Proposition 3.1.
@@
Let the fee rate be constant, that is, $c(t,x)=\bar{c}>0$, and let $g(t,x)=e^{-k(T-t)}$} for some $k>0$. To eliminate the incentive to surrender, it suffices to choose $k$ and $\bar{c}$ such that (9) holds; a quick calculation yields $k\geq\bar{c}$. In this setting, it suffices to set $k$ at or above the fee rate $\bar{c}$ to eliminate optimal early surrenders. This result is well-known and discussed in Bernard and MacKay (2015), Proposition 3.1 and MacKay (2014), Proposition 4.4.2. In particular, MacKay (2014) shows that in the Black-Scholes setting, $g(t,x)=e^{-\bar{c}(T-t)}$ is the minimal surrender charge function that can be used to remove the surrender incentive when the fee rate is constant.
@@
Let $g(t,x)=1-k\left(1-\frac{t}{T}\right)^{3}$ for some constant $k>0$, as in Bacinello and Zoccolan (2019), MacKay (2014) Section 4.3.2, and MacKay et al. (2017). As explained in MacKay (2014), this form of surrender charge function mimics the surrender charges in the market, which are usually high in the first years of the contract and drop drastically thereafter. Assume also that the fee function depends only on time such that $c(t,x)=c(t)$ for all $x\in\mathbb{R}_{+}$. In this example, we do not specify a particular form for the fee function; instead, (9) is used to obtain the form $c(t)$ eliminating the surrender incentive. Simple calculations show that (9) is satisfied if
@@
Thus, any function $c(t)$ that satisfies this inequality will eliminate surrender incentives. For example, one could simply set $c(t)=\frac{3k/T\left(1-t/T\right)^{2}}{1-k(1-t/T)^{3}}$. This example highlights the interplay between the fee and the surrender charge structures. In this setting, it would have been impossible to satisfy (9) using a constant fee rate, except with $c=0$, as the function on the right-hand side of (10) vanishes at maturity. Having a zero fee rate throughout the life of the VA contract is not a feasible solution from the insurer’s perspective since the maturity guarantee must be financed. On the other hand, isolating $k$ in (10) gives
@@
The examples above show that in order to efficiently eliminate early surrender incentives, surrender charges and fees should be structured conjointly.
@@
3.2. General Case
In this section, we study the optimal surrender strategy for the variable annuity contract in the general case, when condition (9) is not necessarily satisfied. We first recall two definitions from probability and optimal stopping theory.